HELP!! I keep getting these two questions wrong.

THANK YOU to whoever can help me with this.

If I dissolve KOH in solution to a concentration of 3.4 x 10-4, what is [OH-] in solution?"

3.4 x 10-4

2.94 x 10-11

Can t be determined without the pKa

2.94 x 10-10

3.47

If [H+] in aqueous solution is 6.50 x 10-6, what is [OH-]?"

1.67 x 10-4

1.67 x 10-9

4.22

1.67 x 10-10

6.50 x 10-6
Answer #1

1) KOH is a strong electrolyte thus dissociate completely and give K+ and OH-. Thus, concentration of OH- = concentration of KOH.

Concentration of OH- = 3.4×10-4

2) [OH-] = 10-14/[H+] = 10-14/6.50×10-6

= 1.67×10-9